# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model
@@
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$
@@
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
@@ B.2 Probit Model
Table B.3: Comparison between the Bias-corrected and Uncorrected MLE
@@
Notes: The true structure parameters are $\theta_{0}=(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $\theta_{0}=\beta_{0}=0.5$ for the static model. In the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ represent the uncorrected MLE and our bias-corrected MLE, respectively. In the dynamic model, $(\widehat{\beta},\widehat{\rho})$ is the uncorrected MLE and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ represents our bias-corrected MLE with the truncation parameter $\tau$ set to $\tau=1$ or $2$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests Based on Bias-corrected and Uncorrected Likelihoods for Probit Model
@@
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model
@@
Notes: The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.

### v4

@@ B.1 Logit Model
Table B.1: Empirical Size of LM and Wald Tests for Logit Model, Design 1
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.2: Empirical Power of LM and Wald Tests for Logit Model using $L(\theta)$, Design 1
@@
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.3: Comparisons of Uncorrected and Corrected Logit Estimates, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Model
$\widehat{\beta}$ $0.454$ $1.201$ $0.100$ $0.132$ $0.060$ $0.071$
$\widehat{\beta}_{L}$ $0.101$ $0.199$ $0.020$ $0.056$ $0.010$ $0.035$
Dynamic Model
$\widehat{\beta}$ $0.527$ $1.583$ $0.109$ $0.212$ $0.059$ $0.071$
$\widehat{\beta}_{L}^{(1)}$ $0.117$ $0.204$ $0.024$ $0.058$ $0.009$ $0.038$
$\widehat{\beta}_{L}^{(2)}$ $0.126$ $0.278$ $0.025$ $0.059$ $0.009$ $0.041$
$\widehat{\rho}$ $-0.110$ $0.241$ $-0.055$ $0.110$ $-0.037$ $0.071$
$\widehat{\rho}_{L}^{(1)}$ $-0.041$ $0.192$ $-0.017$ $0.090$ $-0.010$ $0.059$
$\widehat{\rho}_{L}^{(2)}$ $-0.035$ $0.196$ $-0.012$ $0.091$ $-0.007$ $0.059$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.4: Empirical Size of LR, LM, and Wald Tests for Logit Model, Design 2
(a) LR Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $47\%$ $20\%$ $43\%$ $18\%$
$(60,60)$ $43\%$ $10\%$ $38\%$ $9\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
(b) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $42\%$ $18\%$ $38\%$ $16\%$
$(60,60)$ $42\%$ $9\%$ $37\%$ $9\%$
$(90,90)$ $41\%$ $9\%$ $34\%$ $7\%$
(c) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $32\%$ $15\%$ $27\%$ $15\%$
$(60,60)$ $41\%$ $9\%$ $36\%$ $8\%$
$(90,90)$ $42\%$ $9\%$ $34\%$ $7\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.5: Empirical Power of LR, LM, and Wald Tests for Logit Model using $L(\theta)$, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $79\%$ $62\%$ $45\%$ $12\%$ $16\%$ $26\%$
$(60,60)$ $100\%$ $95\%$ $73\%$ $40\%$ $75\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $75\%$ $58\%$ $41\%$ $19\%$ $26\%$ $35\%$
$(60,60)$ $100\%$ $94\%$ $69\%$ $44\%$ $80\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $77\%$ $60\%$ $42\%$ $12\%$ $17\%$ $28\%$
$(60,60)$ $99\%$ $95\%$ $72\%$ $41\%$ $77\%$ $93\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $82\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $73\%$ $56\%$ $38\%$ $19\%$ $27\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $80\%$ $94\%$
$(90,90)$ $100\%$ $100\%$ $93\%$ $86\%$ $99\%$ $99\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $75\%$ $57\%$ $39\%$ $10\%$ $17\%$ $29\%$
$(60,60)$ $100\%$ $95\%$ $72\%$ $41\%$ $77\%$ $95\%$
$(90,90)$ $100\%$ $100\%$ $94\%$ $81\%$ $98\%$ $100\%$
Dynamic Model
$(30,30)$ $69\%$ $50\%$ $34\%$ $18\%$ $26\%$ $36\%$
$(60,60)$ $100\%$ $93\%$ $68\%$ $45\%$ $81\%$ $97\%$
$(90,90)$ $100\%$ $100\%$ $92\%$ $86\%$ $99\%$ $100\%$
[…]
Notes: Logit model with $\varepsilon_{it}$ standard logistic. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
@@ B.2 Probit Model
Table B.6: Comparisons of Uncorrected and Corrected Probit Estimates, Design 1
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. For the static model, $\widehat{\beta}$ and $\widehat{\beta}_{L}$ are, respectively, the uncorrected and corrected estimates of $\beta_{0}=0.5$. For the dynamic model $(\widehat{\beta},\widehat{\rho})^{\prime}$ is the uncorrected estimate of $(\rho_{0},\beta_{0})^{\prime}=(0.5,0.5)^{\prime}$ from and $(\widehat{\beta}_{L}^{(\tau)},\widehat{\rho}_{L}^{(\tau)})^{\prime}$ is the corrected estimate with truncation parameter $\tau\in(1,2)$. The number of replications is $1000$.
Table B.7: Empirical Size of LR, LM, and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 1
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.8: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 1
@@
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 1: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,0.01)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},0.25)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},0.25)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
Table B.9: Empirical Size of LM and Wald Tests Based on Corrected and Uncorrected Likelihoods for Probit Model, Design 2
(a) LM Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $69\%$ $34\%$ $59\%$ $30\%$
$(60,60)$ $74\%$ $18\%$ $62\%$ $17\%$
$(90,90)$ $71\%$ $14\%$ $60\%$ $13\%$
(b) Wald Tests
Static Model Dynamic Model
$(N,T)$ Uncorrected Corrected Uncorrected Corrected
$(30,30)$ $37\%$ $29\%$ $25\%$ $26\%$
$(60,60)$ $71\%$ $18\%$ $56\%$ $16\%$
$(90,90)$ $72\%$ $14\%$ $60\%$ $13\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5,0.5)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5$ for the static model, with a significance level $5\%$. The number of replications is $1000$.
Table B.10: Empirical Power of LR, LM, and Wald Tests Based on Bias-corrected Likelihoods for Probit Model, Design 2
(a) LR Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $97\%$ $90\%$ $77\%$ $14\%$ $13\%$ $15\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $34\%$ $70\%$ $83\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $85\%$ $98\%$ $96\%$
Dynamic Model
$(30,30)$ $95\%$ $86\%$ $74\%$ $25\%$ $28\%$ $38\%$
$(60,60)$ $100\%$ $100\%$ $93\%$ $47\%$ $84\%$ $96\%$
$(90,90)$ $100\%$ $100\%$ $100\%$ $90\%$ $99\%$ $100\%$
(b) LM Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $96\%$ $87\%$ $73\%$ $12\%$ $13\%$ $17\%$
$(60,60)$ $100\%$ $100\%$ $96\%$ $37\%$ $69\%$ $76\%$
$(90,90)$ $100\%$ $100\%$ $99\%$ $86\%$ $91\%$ $88\%$
Dynamic Model
$(30,30)$ $94\%$ $83\%$ $70\%$ $21\%$ $27\%$ $38\%$
$(60,60)$ $100\%$ $99\%$ $92\%$ $48\%$ $81\%$ $91\%$
$(90,90)$ $100\%$ $99\%$ $99\%$ $90\%$ $94\%$ $94\%$
(c) Wald Tests
$\delta$
$(N,T)$ $-0.20$ $-0.15$ $-0.10$ $0.10$ $0.15$ $0.20$
Static Model
$(30,30)$ $91\%$ $83\%$ $68\%$ $10\%$ $13\%$ $19\%$
$(60,60)$ $99\%$ $99\%$ $95\%$ $39\%$ $76\%$ $92\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $88\%$ $99\%$ $99\%$
Dynamic Model
$(30,30)$ $88\%$ $75\%$ $60\%$ $21\%$ $27\%$ $39\%$
$(60,60)$ $100\%$ $99\%$ $91\%$ $49\%$ $85\%$ $98\%$
$(90,90)$ $99\%$ $99\%$ $99\%$ $91\%$ $99\%$ $100\%$
[…]
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.